how do you find the absolute value of this complex number: 3-6i
Answer by stanbon(75887)

how do you find the absolute value of this complex number: 3-6i
----
|3-6i\ = sqrt[3^2+6^2] = sqrt[9+36]
---
= sqrt(45)
= 3sqrt(5)
===============
Cheers,
Stan H.